OpenStudy (anonymous):

What is the slope of the line 2x – 5y = 12 ?

OpenStudy (anonymous):

rearranging -5y = -2x + 12 y = (2/5)x - 12/5 slope = 2/5

OpenStudy (anonymous):

thanks.